OpenStudy (anonymous):

7% as a decimal and fraction 9.3% as a decimal and fraction 131% as a decimal and fraction i dont understand to divide by 100?

OpenStudy (anonymous):

7% = 0.07 and 7/100 9.3% = 0.093 and 9.3/100 131% = 1.31 and 1 31/100 Since it is a percent all you really have to do is put that number over 100, and then reduce if you can.

OpenStudy (rosedewittbukater):

7% is equal to .07 and 7/100. A percentage is out of a hundred. If you ever get 100% on a test, that means you got 100 out of 100 points. 7% is really 7 out of a hundred, so you divide by 100, or move the decimal point to the left twice. 9.3% would be 0.093 and 9.3/100. Since they probably don't want a decimal in the numerator, you can multiply the top and bottom of the fraction by 10 to move the decimal to the right and get 93/1000. There is a decimal in 100 and all whole numbers, you just can't see them. They're at the very end. So "100." is the same as "100" 131% would be 1.31 because you only move the decimal to the left twice. As a fraction, it would be 131/100. If they want you to simplify so it's not an improper fraction(when the number on top is bigger than the number on the bottom), then make it 1 and 31/100. Hope this helps.

OpenStudy (anonymous):

i have 9.3 over 1000

OpenStudy (anonymous):

okay thanks guys i get it now

OpenStudy (rosedewittbukater):

If you want to multiply the denominator by 10 to make it 1000, you also have to multiply the numerator(the top) to make it 93/1000. Otherwise it's just 9.3/100

OpenStudy (anonymous):

I'm glad you get it. :-)

OpenStudy (anonymous):

okay now like this one i dont get 18% of 50.

OpenStudy (anonymous):

do you just divde them

OpenStudy (rosedewittbukater):

If the question is asking what 18% of 50 is, then first you would have to convert 18% into a decimal like we did before. So divide 18 by 100. Then multiply that number by 50.

OpenStudy (rosedewittbukater):

You only divide by 100 when converting the percentage into a decimal or fraction.

OpenStudy (anonymous):

0.18 divided by 50 =0.0036?

OpenStudy (rosedewittbukater):

0.18 is 18% as a decimal. Since you want to find 18% of 50, multiply 0.18 by 50.

OpenStudy (anonymous):

so 9 percent

OpenStudy (rosedewittbukater):

Just 9. The only percentage is 18, but we changed that to 0.18.

OpenStudy (anonymous):

okay i think im getting it lol thank you :)

OpenStudy (rosedewittbukater):

Glad to help :)

OpenStudy (anonymous):

okay so this one, what % of 148 is 37

OpenStudy (rosedewittbukater):

For this one you would divide 148 by 37. So 148/37 = ? Do you know how to convert that back into a percentage? (hint: you multiply by 100 to turn a percentage into a decimal, then you must do the opposite to turn it back into a percentage)